# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v3

@@ Appendix B Relation with the result in [21]
Similar to this paper, [21] also attempts to model the dynamic of analog training. They shows that Analog SGD converges to a critical point of problem (1) inexactly with an asymptotic error, and Tiki-Taka converges to a critical point exactly. In this section, we compare our results with our results and theirs.
@@
It can be checked that the symmetric point is 0 while the dynamic range of it is $\tau,\tau]$. The symmetric and asymmetric components is defined by $F(W)=1$ and $G(W)=\frac{W}{\tau}$, respectively. It indicates $F_{\max}=1$. Furthermore, they assume the bounded weight saturation by assuming bounded weights, i.e., $\|W_{k}\|_{\infty}\leq W_{\max},\forall k\in[K]$ with a constant $W_{\max}<\tau$. Under this assumption, the lower bounds of response functions are given by
@@
which is a weighted average of $P_{k}$ and $\nabla f(\bar{W}_{k};\xi_{k})$. Consequently, $P_{k}$ can be interpreted as an approximation of the average gradient. For this perspective, the transfer operation can be interpreted as biased gradient descent. However, given a generic $G(\cdot)$, the combination is no longer viable, bringing difficulties to the analysis.
@@
Convergence of Tiki-Taka. It is shown that a non-zero $\gamma$ in (10) improves the training accuracy [22]. However, [21] considers a special case with $\gamma=0$ while this paper considers a non-zero $\gamma$. As we will discuss latter in this section, different $\gamma$ leads to different convergence behaviors of Tiki-Taka.
@@
As a comparison, without introducing a non-zero $\gamma$, [21] shows that convergence rate of Tiki-Taka is only
@@
Even though it is not a completely fair comparison since two paper relies on different assumptions, it is still worthy to compare the analysis in two paper. [21] assumes the noise should be non-zero, i.e. ${\mathbb{E}}_{\xi}[|\nabla f(W;\xi)|]]_{i}\geq c_{\texttt{noise}}\sigma,\forall i\in[D]$ holds for a non-zero constant $c_{\texttt{noise}}$. Instead, this paper does not make this assumption but assumes that the objective is strongly convex. As mentioned in Section 4, the strong convexity is introduced only to ensure the existence of $P^{*}(W_{k})$. Therefore, we believe it can be relaxed, and the convergence rate can remain unchanged, which is left as a future work. Taking that into account, we believe the comparison can provide insight of how the non-zero $\gamma$ improves the convergence rate of Tiki-Taka.
@@
Instead, without a non-zero $\gamma$, [21] points out that $P_{k}$ is an approximation of clear gradient by showing
@@
where $C,C^{\prime}$ are constants depending on the resistive element and model dimension, and the “remainder” is the non-essential terms. Consider the case that $W_{k}$ is fixed and (10) is kept iterating, in which case the increment on $P_{k}$ is constant since $\gamma=0$. Telescoping (23) we find the upper bound above only guarantee that
@@
which means that $P_{k}$ tracks the gradient accurately only when $\nabla f(W_{k})$ reaches zero asymptotically. Consequently, the less accurate approximation leads to a slower rate than this paper.
Appendix C Dynamic of Non-ideal Analog Update
This section presents details about how to get the dynamic of analog update (3) appearing in Section˜2 and its error analysis. The primary distinction between digital and analog training is the weight update method. As discussed in Section 1, the weight update in AIMC hardware is implemented by Analog Update, which sends a series of pulses to the resistive elements.
@@ Appendix C Dynamic of Non-ideal Analog Update
Roughly speaking, given an ideal response $q_{+}(w)=q_{-}(w)=1$ and $\sigma_{c}=0$, $\operatorname{BL}$ pulses, with $\Delta w_{\min}$ increment for each individual pulse, incur the weight update $\Delta w$. Since the response granularity $\Delta w_{\min}$ is scaled by the response function $q_{s}(w)$, the expected increment is approximately scaled by $q_{s}(w)$ as well. Accordingly, we propose an approximate dynamic of Analog Update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined in (3). The following theorem provides an estimation of the approximation error. It has been shown empirically that the response granularity can be made sufficiently small for updating [81, 82], implying $\Delta w_{\min}\ll\Delta w$. Therefore, we establish the error estimation of the approximation based on a small response granularity condition.
@@
The proof of Theorem 4 is deferred to Appendix F. In Theorem 4, $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-U_{q}(w,\Delta w)|$ is the error between the true update and the proposed dynamic, while $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-w|$ is the difference between original weight and the updated one. Theorem 4 shows that the proposed dynamic dominates the update, and the approximation error is negligible when $\Delta w$ is small, which holds as $\Delta w$ always includes a small learning rate in gradient-based training.
Takeaway. Theorem 4 enables us to discuss the impact of response functions directly without dealing with element-specific details like response granularity $\Delta w_{\min}$ and cycle variation $\sigma_{c}$. Response functions are the bridge between the resistive element level equation (pulse update (25)) and algorithm level equation (dynamic of Analog Update (3)).
@@ Appendix D Comparison of Residual Learning v2 and Tiki-Taka v2
which do not include a decay coefficient in front of $H_{k}$. Furthermore, Tiki-Taka v2 uses the gradient $\nabla f(W_{k};\xi_{k})$ that are solely computed on the main array $W_{k}$. Instead, Residual Learning v2 computes gradient on a mixed weight $\bar{W}_{k}=W_{k}+\gamma P_{k}$. As suggested by the ablation simulations in Section˜K.7, the training benefits from a non-zero $\gamma$.
Appendix E Useful Lemmas and Proofs
E.1 Lemma 1: Properties of weighted norm
@@
E.2 Lemma 2: Properties of weighted norm
@@
E.3 Lemma 3: Lipschitz continuity of analog update
@@
E.4 Lemma 4: Element-wise product error
@@
$C_{+}:=$ $\frac{1}{2}\left(\max\{Q\}+\min\{Q\}\right),$ (33)
$C_{-}:=$ $\frac{1}{2}\left(\max\{Q\}-\min\{Q\}\right).$ (34)
@@
$\ \left\langle U,V\odot Q\right\rangle=\sum_{i\in[D]}[U]_{i}[V]_{i}[Q]_{i}$ (35)
$=$ $\ \sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}[U]_{i}[V]_{i}[Q]_{i}\quad+\sum_{i\in[D],[U]_{i}[V]_{i}<0}[U]_{i}[V]_{i}[Q]_{i}$
$\geq$ $\ \min\{Q\}\times\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}[U]_{i}[V]_{i}\right)+\max\{Q\}\times\left(\sum_{i\in[D],[U]_{i}[V]_{i}<0}[U]_{i}[V]_{i}\right)$
$\overset{(a)}{=}$ $\ C_{+}\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}[U]_{i}[V]_{i}\right)-C_{-}\left(\sum_{i\in[D],[U]_{i}[V]_{i}\geq 0}|[U]_{i}[V]_{i}|\right)$
$+$ $\ C_{+}\left(\sum_{i\in[D],[U]_{i}[V]_{i}<0}[U]_{i}[V]_{i}\right)-C_{-}\left(\sum_{i\in[D],[U]_{i}[V]_{i}<0}|[U]_{i}[V]_{i}|\right)$
$=$ $\ C_{+}\sum_{i\in[D]}[U]_{i}[V]_{i}-C_{-}\sum_{i\in[D]}|[U]_{i}[V]_{i}|$
@@
Appendix F Proof of Theorem 4: Error from Pulse Update
See 4
Proof of Theorem 4.
Recall the definition of the bit length is
leading to
@@ Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
$[R_{c}(W)]_{i}:=\int_{\tau_{i}^{\min}}^{[W]_{i}}~[R(W)]_{i}~{\mathrm{d}[W]_{i}}$ (41)
which satisfies $\nabla R_{c}(W)=R(W)$ and $\nabla\left\langle\Sigma,R_{c}(W)\right\rangle=\Sigma\odot R(W)$. Since we do not further assume stronger properties for response functions, like monotonicity, it is hard to provide strong claims on the shape of $R(W)$ or $R_{c}(W)$. Here we provide the expression of $R_{c}(W)$ for the linear response functions $Q_{+}(W)=1-\frac{W}{\tau}$, $Q_{-}(W)=1+\frac{W}{\tau}$. In this case, $F(W)=1$ and $G(W)=\frac{W}{\tau}$ based on definition (4); and hence $R(W)=\frac{G(W)}{F(W)}=\frac{W}{\tau}$. Accordingly, the accumulated asymmetric function is given by
$[R_{c}(W)]_{i}=$ $\ \int_{\tau_{i}^{\min}}^{[W]_{i}}~[R(W)]_{i}~{\mathrm{d}[W]_{i}}=\int_{\tau_{i}^{\min}}^{[W]_{i}}~\frac{[W]_{i}}{\tau}~{\mathrm{d}[W]_{i}}$ (42)
$=$ $\ \frac{1}{2\tau}([W]_{i})^{2}-\frac{1}{2\tau}(\tau_{i}^{\min})^{2}.$
@@
$\left\langle\Sigma,R_{c}(W)\right\rangle=$ $\ \sum_{i=1}^{D}[\Sigma]_{i}[R_{c}(W)]_{i}=\sum_{i=1}^{D}[\Sigma]_{i}\left(\frac{1}{2\tau}([W]_{i})^{2}-\frac{1}{2\tau}(\tau_{i}^{\min})^{2}\right)$ (43)
@@
which is a weighted $\ell_{2}$ norm regularization term. In the scalar case, it reduces to (44).
$\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ (44)
It is noticeable that if the ratio $R(W)$ is monotonic at each coordinate, $R_{c}(W)$ reaches its minimum at $W^{\diamond}$. Therefore, $R_{c}(W)$ has no impact on the convergence only when the optimal point of $f(W)$ is $W^{\diamond}$.
@@ Appendix H Proof of Theorem 2: Convergence of Analog SGD
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (64)
@@
$E_{K}=$ $\ \frac{1}{K}\sum_{k=0}^{K}{\mathbb{E}}[\|\nabla f(W_{k})\odot F(W_{k})+|\nabla f(W_{k})|\odot G(W_{k})\|^{2}+4\|\nabla f(W_{k})\|^{2}_{R(W_{k})}]$ (68)
@@ I.1 Main proof
Let $C=\frac{10\beta F_{\max}^{2}}{\alpha R^{\texttt{RL}}_{\min}\gamma}$, which leads to the positive coefficient in front of $\|P_{k+1}-P^{*}(W_{k})\|^{2}$, i.e.

### v4

@@ Appendix B Relation with the result in [21]
Similar to this paper, [21] also attempts to model the dynamics of analog training. They show that Analog SGD converges to a critical point of problem (1) inexactly with an asymptotic error, and Tiki-Taka converges to a critical point exactly. In this section, we compare our results with our results and theirs.
@@
It can be checked that the symmetric point is 0 while the dynamic range of it is $\tau,\tau]$. The symmetric and asymmetric components are defined by $F(W)=1$ and $G(W)=\frac{W}{\tau}$, respectively. It indicates $F_{\max}=1$. Furthermore, they assume the bounded weight saturation by assuming bounded weights, i.e., $\|W_{k}\|_{\infty}\leq W_{\max},\forall k\in[K]$ with a constant $W_{\max}<\tau$. Under this assumption, the lower bounds of response functions are given by
@@
which is a weighted average of $P_{k}$ and $\nabla f(\bar{W}_{k};\xi_{k})$. Consequently, $P_{k}$ can be interpreted as an approximation of the average gradient. From this perspective, the transfer operation can be interpreted as biased gradient descent. However, given a generic $G(\,\cdot\,)$, the combination is no longer viable, bringing difficulties to the analysis.
@@
Convergence of Tiki-Taka. It is shown empirically that a non-zero $\gamma$ in (10) improves the training accuracy [22]. However, [21] only considers $\gamma=0$ while this paper considers a non-zero $\gamma$.
@@
As a comparison, without a non-zero $\gamma$, [21] shows that convergence rate of Tiki-Taka is only
@@
Even though it is not a completely fair comparison, since the two papers rely on different assumptions, it is still worth comparing their analyses. [21] assumes the noise should be non-zero, i.e. ${\mathbb{E}}_{\xi}[|\nabla f(W;\xi)|]]_{d}\geq c_{\texttt{noise}}\sigma,\forall d\in[D]$ holds for a non-zero constant $c_{\texttt{noise}}$. Instead, this paper does not make this assumption but assumes that the objective is strongly convex. As mentioned in Section 4, the strong convexity is introduced only to ensure the existence of $P^{*}(W_{k})$. Therefore, we believe it can be relaxed and that the convergence rate can remain unchanged, which is left for future work. Taking that into account, we believe the comparison can provide insight into how the non-zero $\gamma$ improves the convergence rate of Tiki-Taka.
@@
Instead, without a non-zero $\gamma$, [21] interprets $P_{k}$ as an approximation of clear gradient by showing
@@
where $C,C^{\prime}$ are constants depending on the resistive element and model dimension, and the “remainder” is the non-essential terms. Consider the case that $W_{k}$ is fixed and (10) is kept iterating, in which case the increment on $P_{k}$ is constant since $\gamma=0$. Telescoping (23), we find that the upper bound above only guarantees that
@@
which means that $P_{k}$ tracks the gradient accurately only when $\nabla f(W_{k})$ reaches zero asymptotically. The less accurate approximation results in a slower rate than the one reported in this paper.
Appendix C Dynamics of Non-ideal Analog Update
This section presents details on how to obtain the dynamics of the analog update (3) appearing in Section˜2, along with its error analysis. The primary distinction between digital and analog training is the method of updating the weight. As discussed in Section 1, the weight update in AIMC hardware is implemented by Analog Update, which sends a series of pulses to the resistive elements.
@@ Appendix C Dynamics of Non-ideal Analog Update
Roughly speaking, given an ideal response $q_{+}(w)=q_{-}(w)=1$ and $\sigma_{c}=0$, $\operatorname{BL}$ pulses, with $\Delta w_{\min}$ increment for each individual pulse, incur the weight update $\Delta w$. Since the response granularity $\Delta w_{\min}$ is scaled by the response function $q_{s}(w)$, the expected increment is approximately scaled by $q_{s}(w)$ as well. Accordingly, we propose an approximate dynamics of Analog Update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined in (3). The following theorem provides an estimation of the approximation error. It has been shown empirically that the response granularity can be made sufficiently small for updating [81, 82], implying $\Delta w_{\min}\ll\Delta w$. Therefore, we establish the error estimate for the approximation under a small-response-granularity condition.
@@
In Theorem 4, $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-U_{q}(w,\Delta w)|$ is the error between the true update and the proposed dynamics, while $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-w|$ is the difference between original weight and the updated one. Theorem 4 shows that the proposed dynamics dominate the update, and the approximation error is negligible when $\Delta w$ is small, which holds as $\Delta w$ always includes a small learning rate in gradient-based training.
Takeaway. Theorem 4 enables us to discuss the impact of response functions directly without dealing with element-specific details like response granularity $\Delta w_{\min}$ and cycle variation $\sigma_{c}$. Response functions are the bridge between the resistive element level equation (pulse update (25)) and the algorithm level equation (dynamics of Analog Update (3)).
Proof of Theorem 4.
Recall the definition of the bit length is
leading to
@@ Appendix D Comparison of Residual Learning v2 and Tiki-Taka v2
which do not include a decay coefficient in front of $H_{k}$. Furthermore, Tiki-Taka v2 uses the gradient $\nabla f(W_{k};\xi_{k})$ that are solely computed on the main array $W_{k}$. Instead, Residual Learning v2 computes gradient on a mixed weight $\bar{W}_{k}=W_{k}+\gamma P_{k}$. As suggested by the ablation simulations in 6, the training benefits from a non-zero $\gamma$.
Appendix E Estimation of time consumption
Residual Learning introduces an extra resistive element array, which increases overhead. However, the extra overhead is affordable in practice. Compared to Analog SGD, the analog memory requirement doubles, but the latency remains almost unchanged since Residual Learning does not explicitly compute the mixed weights during the forward and backward passes. As [83] suggests, $W_{k}$ and $P_{k}$ can share the same analog-digital convertor (ADC), which implements the weight mixing without introducing extra latency. On the other hand, as suggested by [22], the forward, backward, and update steps for $W_{k}$ and $P_{k}$ are performed in parallel, thereby avoiding a significant increase in latency. Consequently, introducing an extra residual array does not incur substantial extra latency.
Following the evaluation in Table 1 in [24], we compared the latency of Analog SGD and Residual Learning in Table˜3. We consider that each gradient update step requires $32$ pulse cycles, each consuming $5$ nanoseconds (ns). Following the estimation in [24], preprocessing the input vectors for each MVM operator takes 5.9ns. Compared with Analog SGD, Residual Learning adds an extra MVM step to read from $P_{k}$. The results suggest that the overhead is only about $2\times$ that of Analog SGD. As the update is typically not the bottleneck of the whole training process, the extra overhead is affordable.
Analog SGD Residual Learning
Forward/backward [84] 40.0 40.0
Update 165.9 371.8
Table 3: Comparison of time (nanosecond) consumption in each layer
Appendix F Useful Lemmas and Proofs
F.1 Lemma 1: Properties of weighted norm
@@
F.2 Lemma 2: Properties of weighted norm
@@
F.3 Lemma 3: Lipschitz continuity of analog update
@@
F.4 Lemma 4: Element-wise product error
@@
$C_{+}:=\frac{1}{2}\left(\max\{Q\}+\min\{Q\}\right),\quad C_{-}:=\frac{1}{2}\left(\max\{Q\}-\min\{Q\}\right).$ (36)
@@
$\ \left\langle U,V\odot Q\right\rangle=\sum_{d\in[D]}[U]_{d}[V]_{d}[Q]_{d}$ (37)
$=$ $\ \sum_{d\in[D],[U]_{d}[V]_{d}\geq 0}[U]_{d}[V]_{d}[Q]_{d}\quad+\sum_{d\in[D],[U]_{d}[V]_{d}<0}[U]_{d}[V]_{d}[Q]_{d}$
$\geq$ $\ \min\{Q\}\times\left(\sum_{d\in[D],[U]_{d}[V]_{d}\geq 0}[U]_{d}[V]_{d}\right)+\max\{Q\}\times\left(\sum_{d\in[D],[U]_{d}[V]_{d}<0}[U]_{d}[V]_{d}\right)$
$\overset{(a)}{=}$ $\ C_{+}\left(\sum_{d\in[D],[U]_{d}[V]_{d}\geq 0}[U]_{d}[V]_{d}\right)-C_{-}\left(\sum_{d\in[D],[U]_{d}[V]_{d}\geq 0}|[U]_{d}[V]_{d}|\right)$
$+$ $\ C_{+}\left(\sum_{d\in[D],[U]_{d}[V]_{d}<0}[U]_{d}[V]_{d}\right)-C_{-}\left(\sum_{d\in[D],[U]_{d}[V]_{d}<0}|[U]_{d}[V]_{d}|\right)$
$=$ $\ C_{+}\sum_{d\in[D]}[U]_{d}[V]_{d}-C_{-}\sum_{d\in[D]}|[U]_{d}[V]_{d}|$
@@ Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
$[R_{c}(W)]_{d}:=\int_{\tau_{i}^{\min}}^{[W]_{d}}~[R(W)]_{d}~{\mathrm{d}[W]_{d}}$ (40)
which satisfies $\nabla\left\langle\Sigma,R_{c}(W)\right\rangle=\Sigma\odot R(W)$ for any $\Sigma\in{\mathbb{R}}^{D}$. Since we do not further assume stronger properties for response functions, like monotonicity, it is hard to provide strong claims on the shape of $R(W)$ or $R_{c}(W)$. Here we provide the expression of $R_{c}(W)$ for the linear response functions $Q_{+}(W)=1-\frac{W}{\tau}$, $Q_{-}(W)=1+\frac{W}{\tau}$. In this case, $F(W)=1$ and $G(W)=\frac{W}{\tau}$ based on definition (4); and hence $R(W)=\frac{G(W)}{F(W)}=\frac{W}{\tau}$. Accordingly, the accumulated asymmetric function is given by
$[R_{c}(W)]_{d}=$ $\ \int_{\tau_{i}^{\min}}^{[W]_{d}}~[R(W)]_{d}~{\mathrm{d}[W]_{d}}=\int_{\tau_{i}^{\min}}^{[W]_{d}}~\frac{[W]_{d}}{\tau}~{\mathrm{d}[W]_{d}}$ (41)
$=$ $\ \frac{1}{2\tau}([W]_{d})^{2}-\frac{1}{2\tau}(\tau_{i}^{\min})^{2}.$
@@
$\left\langle\Sigma,R_{c}(W)\right\rangle=$ $\ \sum_{i=1}^{D}[\Sigma]_{d}[R_{c}(W)]_{d}=\sum_{i=1}^{D}[\Sigma]_{d}\left(\frac{1}{2\tau}([W]_{d})^{2}-\frac{1}{2\tau}(\tau_{i}^{\min})^{2}\right)$ (42)
@@
which is a weighted $\ell_{2}$ norm regularization term. In the scalar case, it reduces to
$\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}.$ (43)
If the ratio $R(W)$ is monotonic at each coordinate, $R_{c}(W)$ reaches its minimum at $W^{\diamond}$. Therefore, $R_{c}(W)$ has no impact on the convergence only when the optimal point of $f(W)$ is $W^{\diamond}$.
@@ Appendix H Proof of Theorem 2: Convergence of Analog SGD
$\ \mathbb{E}_{\xi_{k}}[f(W_{k+1})]$ (63)
$\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
$E^{\texttt{ASGD}}_{K}=$ $\ \frac{1}{K}\sum_{k=0}^{K}{\mathbb{E}}[\|\nabla f(W_{k})\odot F(W_{k})+|\nabla f(W_{k})|\odot G(W_{k})\|^{2}+4\|\nabla f(W_{k})\|^{2}_{R(W_{k})}]$ (67)
@@ I.1 Main proof
Let $C=\frac{10\beta F_{\max}^{2}}{\alpha R^{\texttt{RL}}_{\min}\gamma}$, which leads to the positive coefficient in front of $\|P_{k+1}-P^{*}(W_{k})\|^{2}$, i.e.,